# MVGenMaster: Scaling Multi-View Generation from Any Image via 3D Priors Enhanced Diffusion Model

### v2

@@ Abstract
We introduce MVGenMaster, a multi-view diffusion model enhanced with 3D priors to address versatile Novel View Synthesis (NVS) tasks. MVGenMaster leverages 3D priors that are warped using metric depth and camera poses, significantly enhancing both generalization and 3D consistency in NVS. Our model features a simple yet effective pipeline that can generate up to 100 novel views conditioned on variable reference views and camera poses with a single forward process. Additionally, we have developed a comprehensive large-scale multi-view image dataset called MvD-1M, comprising up to 1.6 million scenes, equipped with well-aligned metric depth to train MVGenMaster. Moreover, we present several training and model modifications to strengthen the model with scaled-up datasets. Extensive evaluations across in- and out-of-domain benchmarks demonstrate the effectiveness of our proposed method and data formulation. Models and codes will be released at https://github.com/ewrfcas/MVGenMaster/.
@@ 3.1 Dataset and Metric Depth Formulation
To avoid collapsing to a naive solution ($r=0$), we force the minimal $r\geq 0.1\frac{std(D_{SfM})}{std(D)}$. With metric depth maps, we formulate 3D priors as warped RGB pixels $I^{warp}$ and CCMs $C^{warp}$ from reference view $i$ to target view $j$ as:
@@
where $C_{i}$ is obtained by unprojecting 2D coordinates into the normalized 3D canonical space, then encoded by Fourier embedding [67]. We integrate 3D priors warped from multiple reference views. In cases of conflict due to occlusion, we retain only warped $I^{warp}$ and $C^{warp}$ with minimal depth in the camera space. Additionally, the contract function [4] based on medium depth eliminates the infinity far depth prediction. Notably, we warp 1:1 RGB pixels instead of 1/8 latent features for two reasons. Firstly, incorporating latent features requires additional VAE encoding [52], which increases training costs by 20%. Second, using uncompressed RGB pixels enhances performance in camera magnifying scenarios, as confirmed by ablation studies. Moreover, warped CCMs are complementary to warped RGB pixels, providing precise information about positioning and occlusion. This is particularly beneficial when the warped RGB pixels may be ambiguous due to significant viewpoint changes.
@@ 3.2 Multi-View Diffusion Model with 3D Priors
where $z_{r}$ and $z_{t}$ represent latent features of reference and target views; $\mathbf{M}$ is a binary mask that distinguishes reference and target features. Unlike CAT3D [20], we incorporate Plücker ray embedding to confirm dense camera presentations. All the aforementioned features are concatenated to form the input for MVGenMaster. RGB pixels, Plücker ray, and masks are maintained in 1:1 scale, while CCMs and latent features are downscaled to 1/8 scale. A full convolutional model encodes and adds them into input features. Ablation studies in Table 5 show that this straightforward incorporation of 3D priors results in significant improvements. The backbone of MVGenMaster is built upon the v-prediction based SD2 [52], following the design of CAT3D with 3D full attention across all views as well as discarding the textual branch and cross-attention modules. MVGenMaster is trained with $N+M=8$, where $N,M$ mean reference and target view numbers respectively. During the first training phase, we set $N=3$ as CAT3D. In the second phase, the reference views are dynamically adjusted to $N=1\sim 3$. MVGenMaster can be generalized to varying $N$ and $M$ by key-rescaling, as outlined in Section 3.3.
@@ 3.3 Key-Rescaling for View Extension
Solution. We propose the key-rescaling, a training-free technique designed to improve the reference guidance in self-attention modules as shown in Figure 4(b). Formally, MVGenMaster adopts 3D full self-attention across all views, while key-rescaling multiplies a constant $\gamma$ to the reference views of key features in all self-attention modules. Therefore, all reference views would achieve higher dot product scores before the softmax operation, thereby allowing the aggregation to focus more on reference guidance and balance the attention dilution. Empirically, we find that $\gamma=1.2$ performs well in most scenarios when $M>25$. As verified in Figure 4 and our ablation studies, MVGenMaster strengthened by key-rescaling exhibits strong generalization across various target views. All experiments of MVGenMaster in this paper are accomplished by a single forward process. Moreover, the proposed key-rescaling can be seamlessly operated to the key features before the attention module, making it compatible with the widely used FlashAttention2 [14].
@@ 4.2 Results of Reconstruction
We evaluate the performance of 3DGS reconstruction [30] based on NVS results in Table 3. For the 2-view scenario, the first and last frames are provided, while NVS methods produce additional 23 frames; there are 25 frames at all. We compare ViewCafter-sparse, CAT3D*, and our method here, where ViewCrafter-sparse has been fine-tuned with the first and last frames. Our MVGenMaster enjoys preferable performance on zero-shot datasets. For the 3-view scenario, CAT3D* and MVGenMaster generate 100 frames to support the 3DGS. Especially, MVSplat [13] fails when there is insufficient view overlap. Notably, MVGenMaster achieves prominent advancements compared to other methods benefiting from good consistency achieved through key-rescaling.
@@ 5 Conclusion
Limitation. Although MVGenMaster is a powerful NVS model, an interesting future work involves modifying the base model design, e.g., replacing the backbone with Diffusion Transformer (DiT) [48] for superior capacity. Besides, unifying both rendering and multi-view generation is a promising way to improve consistency and stability of NVS.
@@ Appendix B More Implementation Details
B.1 Multi-Scale Training
@@
Resolution (height x width)
320x768 384x448 576x448
320x704 448x576 576x384
320x640 448x512 576x320
320x576 512x512 640x384
320x512 448x384 640x320
384x640 512x448 704x320
384x576 512x384 768x320
384x512 512x320
@@
B.2 Normalization of Camera Poses
@@
B.3 Settings of 3DGS Reconstruction
We follow the vanilla 3DGS implementation of [30] with L1, SSIM, and LPIPS [84] losses as many other works [45, 41, 82]. The dynamic LPIPS weighting strategy [41, 20] is also incorporated according to the distance to the nearest reference view, and the LPIPS weight changes from 0 (reference view) to 0.25 (the farthest view). The training is accomplished within 2000 steps in our work.
B.4 Efficiency and Memory Costs
Our model enjoys a similar efficiency as CAT3D [20], while the cost of 3D priors incorporation is negligible. Inference efficiency and GPU memory costs of the 50-step DDIM scheduler with different target view numbers are listed in Table 9. The experimental device is one 80GB A800 NVIDIA GPU; the image resolution is 384x576. Note that our model can synthesize all views at once instead of cumbersome iterative generations.
@@ Appendix C Supplementary Experiment Results
Limitations. In Figure 10, we show some ambiguous artifacts in the background generated by our model. The unclear and ambiguous background regions of the reference view mainly cause such artifacts. Improving the model backbone with superior capacity would refine this issue as mentioned in our main paper.
Appendix D Broader Impacts

### v3

@@ Abstract
We introduce MVGenMaster, a multi-view diffusion model enhanced with 3D priors to address versatile Novel View Synthesis (NVS) tasks. MVGenMaster leverages 3D priors that are warped using metric depth and camera poses, significantly enhancing both generalization and 3D consistency in NVS. Our model features a simple yet effective pipeline that can generate up to 100 novel views conditioned on variable reference views and camera poses with a single forward process. Additionally, we have developed a comprehensive large-scale multi-view image dataset called MvD-1M, comprising up to 1.6 million scenes, equipped with well-aligned metric depth to train MVGenMaster. Moreover, we present several training and model modifications to strengthen the model with scaled-up datasets. Extensive evaluations across in- and out-of-domain benchmarks demonstrate the effectiveness of our proposed method and data formulation.
@@ 3.1 Dataset and Metric Depth Formulation
To avoid collapsing to a naive solution ($r=0$), we force the minimal $r\geq 0.1\frac{std(D_{SfM})}{std(D)}$. With metric depth maps, we formulate 3D priors as warped RGB pixels $I^{warp}$ and CCMs [36] $C^{warp}$ from reference view $i$ to target view $j$ as:
@@
where the 3D coordinate map $C_{i}$ is obtained by unprojecting 2D coordinates into the normalized 3D canonical space through aligned depth and camera, then encoded by Fourier embedding [69]. We integrate 3D priors warped from multiple reference views. In cases of conflict due to occlusion, we retain only warped $I^{warp}$ and $C^{warp}$ with minimal depth in the camera space. Additionally, the contract function [4] based on medium depth eliminates the infinity far depth prediction. Notably, we warp 1:1 RGB pixels instead of 1/8 latent features for two reasons. Firstly, incorporating latent features requires additional VAE encoding [54], which increases training costs by 20%. Second, using uncompressed RGB pixels enhances performance in camera magnifying scenarios, as confirmed by ablation studies. Moreover, warped CCMs are complementary to warped RGB pixels, providing precise information about positioning and occlusion. This is particularly beneficial when the warped RGB pixels may be ambiguous due to significant viewpoint changes.
@@ 3.2 Multi-View Diffusion Model with 3D Priors
where $z_{r}$ and $z_{t}$ represent latent features of reference and target views; $\mathbf{M}$ is a binary mask that distinguishes reference and target features. Unlike CAT3D [21], we incorporate Plücker ray embedding to confirm dense camera presentations. All the aforementioned features form the input for MVGenMaster. RGB pixels, Plücker ray, and masks are maintained in 1:1 scale, while CCMs and latent features are downscaled to 1/8 scale. Note that we only use VAE to get latent features, while other inputs are encoded by a lightweight CNN trained from scratch and added to latent features. Ablation studies in Table 5 show that this straightforward incorporation of 3D priors results in significant improvements. The backbone of MVGenMaster is built upon the v-prediction based SD2 [54], following the design of CAT3D with 3D full attention across all views as well as discarding the textual branch and cross-attention modules. MVGenMaster is trained with $N+M=8$, where $N,M$ mean reference and target view numbers respectively. During the first training phase, we set $N=3$ as CAT3D. In the second phase, the reference views are dynamically adjusted to $N=1\sim 3$. MVGenMaster can be generalized to varying $N$ and $M$ by key-rescaling, as outlined in Section 3.3.
@@ 3.3 Key-Rescaling for View Extension
Solution. We propose the key-rescaling, a training-free technique to improve the reference guidance in self-attention modules as shown in Figure 4(b). Formally, MVGenMaster adopts 3D full self-attention across all views, while key-rescaling multiplies a constant $\gamma$ to the reference views of key features in all self-attention modules. Therefore, all reference views would achieve higher dot product scores before the softmax operation, thereby allowing the aggregation to focus more on reference guidance and balance the attention dilution. Empirically, we find that $\gamma=1.2$ performs well in most scenarios when 158 $\geq(N+M)>$ 28 (158-view NVS is the upper bound of an 80G GPU). More detailed ablation studies are discussed in the supplementary. As verified in Figure 4 and our ablation studies, MVGenMaster strengthened by key-rescaling exhibits strong generalization across various target views. All experiments of MVGenMaster in this paper are accomplished by a single forward process. Moreover, the proposed key-rescaling can be seamlessly operated to the key features before the attention module, making it compatible with the widely used FlashAttention2 [15].
@@ 4.1 Results of MVGenMaster
Reconstruction. We evaluate the performance of 3DGS reconstruction [32] based on NVS results in Table 3. For the 2-view scenario, the first and last frames are provided, while NVS methods produce other 23 frames; there are 25 frames at all. We compare ViewCafter-sparse, CAT3D*, and our method here, where ViewCrafter-sparse has been fine-tuned with the first and last frames. Our MVGenMaster enjoys preferable performance on zero-shot datasets. For the 3-view scenario, CAT3D* and MVGenMaster generate 100 frames to support the 3DGS. Especially, MVSplat [14] fails when there is insufficient view overlap. Notably, MVGenMaster achieves prominent advancements compared to other methods benefiting from good consistency achieved through key-rescaling. We claim that results from Table 3 are under a challenging but efficient setting, i.e., 3DGS based on once NVS inference, without cumbersome iterative generation [21] and distillation [75]. We further compare our method to other closed-source methods in the supplementary.
@@ Appendix B More Implementation Details
B.1 Canonical Coordinate Map (CCM)
We provide more details about CCM [36] mentioned in the main paper. To incorporate precise positional information to MVGenMaster, we begin by unprojecting 2D coordinates $x_{r}$ from reference views into the 3D world space (canonical space) as:
$C^{world}_{r}\simeq P^{c\rightarrow w}_{r}\hat{D}_{r}(x_{r})K^{-1}_{r}x_{r},$ (5)
where $C^{world}_{r}$ denotes the 3D coordinates in the world space (CCM); $P^{c\rightarrow w}_{r}$ is the transformation matrix converting coordinates from camera to world space; $\hat{D}_{r}(x_{r})$ and $K^{-1}_{r}$ indicate metric depth and camera intrinsic’s inversion of reference views, respectively. Subsequently, $C^{world}_{r}$ is further warped into target views, i.e., $C^{warp}_{r\rightarrow t}$ following Eq. 3 in the main paper. Note that we further normalize $C^{warp}_{r\rightarrow t}$ into 0 to 1 before send it to the Fourier embedding.
B.2 Multi-Scale Training
@@
Resolution (height $\times$ width)
320$\times$768 384$\times$448 576$\times$448
320$\times$704 448$\times$576 576$\times$384
320$\times$640 448$\times$512 576$\times$320
320$\times$576 512$\times$512 640$\times$384
320$\times$512 448$\times$384 640$\times$320
384$\times$640 512$\times$448 704$\times$320
384$\times$576 512$\times$384 768$\times$320
384$\times$512 512$\times$320
@@
B.3 Normalization of Camera Poses
@@
B.4 Settings of 3DGS Reconstruction
We follow the vanilla 3DGS implementation of [32] with L1, SSIM, and LPIPS [86] losses as many other works [47, 43, 84]. The point cloud from Dust3R [72] is leveraged to initialize the 3DGS training. The dynamic LPIPS weighting strategy [43, 21] is also incorporated according to the distance to the nearest reference view, and the LPIPS weight changes from 0 (reference view) to 0.25 (the farthest view). The training is accomplished within 2000 steps in our work.
B.5 Efficiency and Memory Costs
Our model enjoys a similar efficiency as CAT3D [21], while the cost of 3D priors incorporation is negligible. Inference efficiency and GPU memory costs of the 50-step DDIM scheduler with different target view numbers are listed in Table 9. The experimental device is one 80GB A800 NVIDIA GPU; the image resolution is 384$\times$576. Note that our model can synthesize all views at once instead of cumbersome iterative generations.
@@ Appendix C Supplementary Experiment Results
3-view 6-view 9-view
[…]
Real10k Recon[75] 25.84 0.144 29.99 0.103 31.82 0.092
CAT3D[21] 26.78 0.132 31.07 0.092 32.20 0.082
Ours 25.15 0.091 27.28 0.067 27.54 0.062
CO3D Recon[75] 19.59 0.398 21.84 0.342 22.95 0.318
CAT3D[21] 20.57 0.351 22.79 0.292 23.58 0.273
Ours 19.99 0.278 21.96 0.214 22.77 0.195
Mip360 Recon[75] 15.50 0.585 16.93 0.544 18.19 0.511
CAT3D[21] 16.62 0.515 17.72 0.482 18.67 0.460
Ours 16.29 0.489 17.51 0.405 18.21 0.354
Table 10: Reconstruction results of the official Reconfusion benchmark. Results of reconfusion [75] and CAT3D [21] are from their papers. Our results are under once NVS inference and 3DGS.
Reconstruction Benchmark. To ensure the effectiveness of MVGenMaster, we compare our method with the official data splits from the benchmark of Reconfusion [75] as in Table 10. Results of Reconfusion [75] and CAT3D [21] are from their papers. Our method enjoys a much simpler pipeline to produce all views with one forward process without iterative generation (CAT3D) and distillation (Reconfusion), only taking 3 minutes for each scene with one A800 (NVS+3DGS). Besides, Reconfusion and CAT3D use ZipNeRF [5], which costs much more training times for each instance. Moreover, we should clarify that the results of MipNeRF-360 are slightly different from Table.3 of the main paper. Because the data splits of Reconfusion are different from ours. They use a heuristic strategy to encourage more reasonable camera placing for the object centric reconstruction [75]. Even under such a challenging setting, our method achieves better human perception (lower LPIPS) and comparable PSNR.
@@
[…]
1/97 1.0 12.53 0.541 2/97 1.0 13.69 0.468
1.1 12.56 0.528 1.1 14.31 0.432
1.2 12.52 0.523 1.2 14.46 0.426
1.3 12.50 0.526 1.3 14.35 0.438
3/50 1.0 14.82 0.386 3/97 1.0 14.75 0.402
1.1 15.47 0.359 A1 15.59 0.360
1.2 15.55 0.358 1.2 15.64 0.358
1.3 15.28 0.371 A2 15.48 0.364
Table 13: Detailed Ablation of Key-rescaling. Following the adaptive $\gamma$ settings based on attention-entropy [31], A1 is set based on the length of $N+M$, while A2 further considers the various sequential lengths across different layers.
Detailed Ablation of Key-rescaling. We clarified that $\gamma=1.2$ defined in the main paper works robustly enough for our method across various $N,M$ (up to upper bound $N+M$=158). We further provide detailed ablation studies and discussions about key-rescaling in Table 13. i) $N=1$ suffers from ambiguous NVS by generating vast number of views (usually set $M\leq 25$ for $N=1$). As exploratory results, key-rescaling’s effect of $N=1$ is a little weaker than $N>1$, but it still achieves the best quality under $\gamma=1.2$ (best LPIPS). ii) Longer sequence enjoys larger $\gamma$, while shorter one requires smaller $\gamma$. Since $N+M=100$ addresses most scenarios, $\gamma=1.2$ is robust enough for our method. iii) We show adaptive $\gamma$ settings based on attention-entropy [31] (A1, A2). They fail to get superior results compared to constant $\gamma$. Further exploring the key-rescaling theory and finding the optimal $\gamma$ is interesting future work. Our work provides this insight to the community and verifies the proposed key-rescaling is generalized enough to address our problems based on Occam’s Razor.
@@
Limitations. In Figure 10, we show some ambiguous artifacts in the background generated by our model. Such artifacts are mainly caused by the unclear and ambiguous background regions of the reference view. Improving the model backbone with superior capacity would refine this issue as mentioned in our main paper.
Appendix D Discussion and Future Work
Although MVGenMaster is a powerful NVS model, an interesting future work involves modifying the base model design, e.g., replacing the backbone with Diffusion Transformer (DiT) [50] for superior capacity. Besides, unifying both rendering and multi-view generation is a promising way to improve consistency and stability of NVS.
Appendix E Broader Impacts